[Brahms-l] Shift report 20050128 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Jan 28 2005 - 23:50:13 EST
Hiro and Bob replaced the Si CAMAC crate that was giving problems.  After inserting two crates and several power supplies, they found a seemingly unique combination that works.

Had a very nice store starting at 17:30.  The MRS is at 45 deg at A1050.
FS still at 8 deg at A1219
Runs Achieved
Run    tr2     tr3
12647  89k    494k 
12648 105k    231k
12651  49k    150k
12653  57k    175k

Beam was dumped at 22:00.  The first attempt at a new store failed.  They appear to be trying again.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Jan 28 23:50:52 2005

This archive was generated by hypermail 2.1.8 : Fri Jan 28 2005 - 23:51:15 EST